# Analytic Personalized Federated Meta-Learning

### v1

@@ 29       Set the local batch size to $B_{l+1}^{(i)}$;
III-D Theoratical Analysis on Computation and Communication
@@
Computational complexity. In the conventional paradigm of FL, the whole training process contains a number of training rounds and each training round requires each client to conduct several repeated visits (i.e. local epochs) on its whole local dataset. Meanwhile, each epoch includes multiple iterations, in which each client samples a batch of data and labels from $\mathcal{D}_{i}$ without replacement and conducts certain computations on the batch of data and labels. Since each DNN layer in FedACnnL converges in one training round which contains one local epoch, FedACnnL only requires $L$ training rounds in total and each training round contains one local epoch. That is greatly smaller than both FedZO and BAFFLE, in which more than 20/40 training rounds (each contains more than one local epoch) are required to achieve a certain model performance empirically on the MNIST [20]/CIFAR-10 [21] dataset using LeNet-5 (5 layers) [22]. Therefore, our analysis focuses on the computations conducted on the batch of data and labels to prove a lower batch-level computational complexity for FedACnnL. In FedACnnL, the batch-level computations contain one forward pass and three matrix multiplications (line 8 and line 10 $\sim$ 14 in the algorithm 1). However, in FedZO and BAFFLE, the batch-level computations incorporate performing a large number of finite difference processes to accurately approximate the gradient. In each finite difference process, two forward passes should be conducted. After removing the computational complexity of one forward pass for each framework, the complexity of the remaining three matrix multiplications in FedACnnL is $\mathcal{O}(B(d_{y}d_{l}+d_{l-1}^{2}+d_{l-1}d_{l}))$ in the arbitrary $l$-th training round. Meanwhile, more than one forward passes exists in FedZO or BAFFLE and each of them has a complexity of $\mathcal{O}(\sum_{l=1}^{L}Bd_{l-1}d_{l})$. Thus, even if FedZO or BAFFLE merely conducts the finite difference process once (which is impractical) in the batch-level computations, FedACnnL can still save a considerable amount of training time. This theoretical conclusion is empirically verified in the section V-B.
@@
Communication complexity. In each training round, each client in FedACnnL uploads the resultant two items (i.e. $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$) to the PS and receive the trained global model weights of the current training layer $\hat{\textbf{W}}_{l}$ from the PS. Notably, $B_{l}^{(i)}$ is too small and we ignore it. Thus, the total communication complexity for FedACnnL in the arbitrary $l$-th training round is $\mathcal{O}(d_{l-1}^{2}+d_{l-1}d_{l}+\theta_{l})$, where $\theta_{l}$ denotes the parameter number of the $l$-th DNN layer. Meanwhile, each client in FedZO or BAFFLE uploads and receives the entire DNN model in each training round, resulting in a communication complexity of $\mathcal{O}(2\sum_{l=1}^{L}\theta_{l})$ per round. Whether in the convolutional layer or the fully-connected layer, $\theta_{l}$ is equal to $d_{l-1}d_{l}$. Therefore, FedACnnL can achieve a lower communication complexity than FedZO and BAFFLE unless the output dimensionality of a certain layer exceeds the parameter number of the entire DNN model. That is a rare case. Furthermore, since FedACnnL requires only $L$ training rounds while FedZO and BAFFLE need a great number one empirically (see section V-B), the total communication complexity of FedACnnL in the whole FL process is much lower than FedZO and BAFFLE.
IV Design of pFedACnnL
IV-A Overview of pFedACnnL
@@
The initialization stage. In the same way as FedACnnL, each client utilizes the random seed $s$ received from the PS to generate an identical set of label encoding matrices $\mathcal{Q}$. Then, a local model is initialized with the same structure on each client. Next, different from FedACnnL, pFedACnnL requires each client to upload an encoded label vector, which is the representation of the local data distribution on each client, to the PS. Finally, the PS uses these encoded label vectors to conduct the $K$-means clustering algorithm and classifies the clients into $K$ groups. In this way, pFedACnnL enables the clients that have similar local data distribution to train a joint meta-model, namely the group-wise meta-model, in the next stage. By doing so, each group-wise meta-model can contain a specialized unique common knowledge within its group and make it easier for its clients to conduct local personalization than a single global model (see section V-C).
@@ IV-B The Initialization Stage
IV-C The Federated Optimization Stage
@@
IV-D The Local Personalization Stage
@@
V Performance Evaluation
V-A Experimental Settings
@@
Figure 5: The overall test accuracy of all clients versus training round cure during the training procedure on the MNIST and synthetic datasets.
V-B Training Efficiency
To highlight the efficiency of FedACnnL on training DNN, we carry out the experiment in the non-convex setting (i.e. DNN) to validate the training time reduction of FedACnnL over the aforementioned first- and zeroth-order FL frameworks (i.e. FedAvg, FedACnnL-WA, FedZO and BAFFLE). Before we discuss the training time, we should establish the model performance guarantee for FedACnnL. Fig. 5 demonstrates the overall test accuracy of all clients versus training round curve during the training procedure. It shows that FedACnnL can approach a competitive model performance with the first-order FL framework FedAvg on the MNIST dataset while it achieves the state-of-the-art (SOTA) model performance on the synthetic dataset. Meanwhile, FedACnnL-WA achieves the same model performance as FedACnnL due to the weight-invariant property which has been discussed in the section III-B. On the basis of the performance superiority, we present the averaged total training time for each framework in Fig. 4. As illustrated in Fig. 4, FedACnnL can reduce 88%, 98% and 99% training time of FedAvg, FedZO and BAFFLE respectively while the reduction ratio is more than 98% on the synthetic dataset. That is because FedACnnL updates its model weights only once in a layer-wise manner while the other frameworks require clients to iteratively update their models locally and need substantial training rounds. Furthermore, the total time per training round in FedZO and BAFFLE is $4.7s,7.0s$ on the MNIST dataset and $0.31s,1.07s$ on the synthetic dataset while that in FedACnnL is $0.42s$ and $0.15s$ on the MNIST and synthetic datasets. This fact validates our theoretical analysis in the section III-D which indicates that FedACnnL has a lower complexity of computation and communication on each client in each training round than FedZO and BAFFLE. Finally, FedACnnL reduces $13\%$ and $21\%$ idle time of FedACnnL-WA on the MNIST and synthetic dataset, proving the effectiveness of the adaptive batch size algorithm in which the time expenses of the local computation for the lagging clients are reduced.
@@ V-B Training Efficiency
TABLE II: Table of model performance.
@@
V-C Personalized Model Performance
We evaluate our proposed APFML framework in both the convex and non-convex settings. The evaluation is done by taking the mean and standard deviation of the test accuracy from running the experiment for 5 times. Table II lists the model performance of pFedACnnL, the vanilla FedACnnL and another four PFL frameworks. In the table, we can find that pFedACnnL experiences a performance improvement of $4\%\sim 6\%$ in the test accuracy under the non-convex setting while it has an improvement of $7\%\sim 8\%$ under the convex setting, compared to the vanilla FedACnnL. Furthermore, as illustrated in Fig. 6 and Table II, in most cases (except in the MNIST dataset), pFedACnnL has a better model than its versions without a certain component (i.e. pFedACnnL-WC and pFedACnnL-WP). For instance, in Fig. 6, pFedACnnL achieves a test accuracy of 74.40% on the CIFAR-10 dataset while pFedACnnL-WC and pFedACnnL-WP achieve a test accuracy of 72.88% and 74.03% respectively. Notably, on the CIFAR-10 and synthetic datasets, compared with pFedACnnL-WP, pFedACnnL-WC which uses a single global model has an even lower model performance than pFedACnnL, proving the effectiveness of the grouping mechanism in the highly heterogeneous data environment. The special case is that pFedACnnL-WC outperforms pFedACnnL in the MNIST dataset. That is because the task difficulty of the MNIST dataset is comparatively lower than that of the synthetic and CIFAR-10 datasets. Thus, introducing too many complex techniques (including DNN, group-wise meta-learning and etc) to pFedACnnL for easy machine learning tasks can cause an over-fitting problem. Last but not least, pFedACnnL achieves the SOTA performance compared to the first- and zeroth-order PFL frameworks (i.e. pFedMe and pFedZO) in most cases while it achieves a competitive performance with pFedMe (96.23% versus 96.92%) in the non-convex setting of the MNIST dataset.
VI Conclusions
In this paper, we have proposed FedACnnL to enable computation-efficient DNN training in AFL and proposed pFedACnnL based on it to mitigate the heterogeneous data distribution problem in AFL. Specifically, FedACnnL is the federated implementation of ACnnL with an introduction of the adaptive batch size algorithm to boost the training procedure while pFedACnnL incorporates a gradient-free meta-learning method on the basis of FedACnnL. Experimentally, FedACnnL has a significantly shorter training time than the conventional first- and zeroth-order FL frameworks while pFedACnnL achieves the SOTA model performance in most cases of convex and non-conv setting.

### v2

@@ 29       Set the local batch size to $B_{l+1}^{(i)}$;
3.4 Theoretical Analysis on Computation and Communication
@@
Computational complexity. In the conventional paradigm of FL, the whole training process contains a number of training rounds and each training round requires each client to conduct several repeated visits (i.e. local epochs) on its whole local dataset. Meanwhile, each epoch includes multiple iterations, in which each client samples a batch of data and labels from $\mathcal{D}_{i}$ without replacement and conducts certain computations on the batch of data and labels. Since each DNN layer in FedACnnL converges in one training round which contains one local epoch, FedACnnL only requires $L$ training rounds in total and each training round contains one local epoch. That is greatly smaller than both FedZO and BAFFLE, in which more than 20/40 training rounds (each contains more than one local epoch) are required to achieve a certain model performance empirically on the MNIST [20]/CIFAR-10 [21] dataset using LeNet-5 (5 layers) [22]. Therefore, our analysis focuses on the computations conducted on the batch of data and labels to prove a lower batch-level computational complexity for FedACnnL. In FedACnnL, the batch-level computations contain one forward pass and three matrix multiplications (line 8 and line 10 $\sim$ 14 in the algorithm 1). However, in FedZO and BAFFLE, the batch-level computations incorporate performing a large number of finite difference processes to accurately approximate the gradient. In each finite difference process, two forward passes should be conducted. After removing the computational complexity of one forward pass for each framework, the complexity of the remaining three matrix multiplications in FedACnnL is $\mathcal{O}(B(d_{y}d_{l}+d_{l-1}^{2}+d_{l-1}d_{l}))$ in the arbitrary $l$-th training round. Meanwhile, more than one forward passes exists in FedZO or BAFFLE and each of them has a complexity of $\mathcal{O}(\sum_{l=1}^{L}Bd_{l-1}d_{l})$. Thus, even if FedZO or BAFFLE merely conducts the finite difference process once (which is impractical) in the batch-level computations, FedACnnL can still save a considerable amount of training time. This theoretical conclusion is experimentally verified in the section 5.2.
@@
Communication complexity. In each training round, each client in FedACnnL uploads the resultant two items (i.e. $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$) to the PS and receive the trained global model weights of the current training layer $\hat{\textbf{W}}_{l}$ from the PS. Notably, $B_{l}^{(i)}$ is too small and we ignore it. Thus, the total communication complexity for FedACnnL in the arbitrary $l$-th training round is $\mathcal{O}(d_{l-1}^{2}+d_{l-1}d_{l}+\theta_{l})$, where $\theta_{l}$ denotes the parameter number of the $l$-th DNN layer. Meanwhile, each client in FedZO or BAFFLE uploads and receives the entire DNN model in each training round, resulting in a communication complexity of $\mathcal{O}(2\sum_{l=1}^{L}\theta_{l})$ per round. Whether in the convolutional layer or the fully-connected layer, $\theta_{l}$ is equal to $d_{l-1}d_{l}$. Therefore, FedACnnL can achieve a lower communication complexity than FedZO and BAFFLE unless the output dimensionality of a certain layer exceeds the parameter number of the entire DNN model. That is a rare case. Furthermore, since FedACnnL requires only $L$ training rounds while FedZO and BAFFLE need a great number one experimentally (see section 5.2), the total communication complexity of FedACnnL in the whole FL process is much lower than FedZO and BAFFLE.
4 Design of pFedACnnL
4.1 Overview of pFedACnnL
@@
The initialization stage. In the same way as FedACnnL, each client utilizes the random seed $s$ received from the PS to generate an identical set of label encoding matrices $\mathcal{Q}$. Then, a local model is initialized with the same structure on each client. Next, different from FedACnnL, pFedACnnL requires each client to upload an encoded label vector, which is the representation of the local data distribution on each client, to the PS. Finally, the PS uses these encoded label vectors to conduct the $K$-means clustering algorithm and classifies the clients into $K$ groups. In this way, pFedACnnL enables the clients that have similar local data distribution to train a joint meta-model, namely the group-wise meta-model, in the next stage. By doing so, each group-wise meta-model can contain a specialized unique common knowledge within its group and make it easier for its clients to conduct local personalization than a single global model (see section 5.3).
@@ 4.2 The Initialization Stage
4.3 The Federated Optimization Stage
@@
4.4 The Local Personalization Stage
@@
5 Performance Evaluation
5.1 Experimental Settings
@@
Figure 5: The overall test accuracy of all clients versus training round curve during the training procedure on the MNIST and synthetic datasets.
5.2 Training Efficiency
To highlight the efficiency of FedACnnL on training DNN, we carry out the experiment in the non-convex setting to validate the training time reduction of FedACnnL over the aforementioned first- and zeroth-order FL frameworks (i.e. FedAvg, FedACnnL-WA, FedZO and BAFFLE). Before we discuss the training time, we should establish the model performance guarantee for FedACnnL. Fig. 5 demonstrates the overall test accuracy of all clients versus training round curve during the training procedure. It shows that FedACnnL can approach a competitive model performance with the first-order FL framework FedAvg on the MNIST dataset while it achieves the state-of-the-art (SOTA) model performance on the synthetic dataset. Meanwhile, FedACnnL-WA achieves the same model performance as FedACnnL due to the weight-invariant property which has been discussed in the section 3.2. On the basis of the performance superiority, we present the averaged total training time for each framework in Fig. 4. As illustrated in Fig. 4, FedACnnL can reduce 83%, 98% and 99% training time of FedAvg, FedZO and BAFFLE respectively while the reduction ratio is more than 97% on the synthetic dataset. That is because FedACnnL updates its model weights only once in a layer-wise manner while the other frameworks require clients to iteratively update their models locally and need substantial training rounds. Furthermore, the total time per training round in FedZO and BAFFLE is $4.7s,7.0s$ on the MNIST dataset and $0.31s,1.07s$ on the synthetic dataset while that in FedACnnL is $0.60s$ and $0.35s$ on the MNIST and synthetic datasets. This fact validates our theoretical analysis in the section 3.4 which indicates that FedACnnL has a lower complexity of computation and communication on each client in each training round than FedZO and BAFFLE. Finally, FedACnnL reduces $52\%$ and $40\%$ idle time of FedACnnL-WA on the MNIST and synthetic dataset, proving the effectiveness of the adaptive batch size algorithm in which the time expenses of the local computation for the lagging clients are reduced.
@@ 5.2 Training Efficiency
Table 2: Table of model performance.
@@
5.3 Personalized Model Performance
We evaluate our proposed APFML framework in both the convex and non-convex settings. The evaluation is done by taking the mean and standard deviation of the test accuracy from running the experiment for 5 times. Table 2 lists the model performance of pFedACnnL, the vanilla FedACnnL and another four PFL frameworks. In the table, we can find that pFedACnnL experiences a performance improvement of $4\%\sim 6\%$ in the test accuracy under the non-convex setting while it has an improvement of $7\%\sim 8\%$ under the convex setting, compared to the vanilla FedACnnL. Furthermore, as illustrated in Fig. 6 and Table 2, in most cases (except in the MNIST dataset), pFedACnnL has a better model than its versions without a certain component (i.e. pFedACnnL-WC and pFedACnnL-WP). For instance, in Fig. 6, pFedACnnL achieves a test accuracy of 74.40% on the CIFAR-10 dataset while pFedACnnL-WC and pFedACnnL-WP achieve a test accuracy of 72.88% and 74.03% respectively. Notably, on the CIFAR-10 and synthetic datasets, compared with pFedACnnL-WP, pFedACnnL-WC which uses a single global model has an even lower model performance than pFedACnnL, proving the effectiveness of the grouping mechanism in the highly heterogeneous data environment. The special case is that pFedACnnL-WC outperforms pFedACnnL in the MNIST dataset. That is because the task difficulty of the MNIST dataset is comparatively lower than that of the synthetic and CIFAR-10 datasets. Thus, introducing too many complex techniques (including DNN, group-wise meta-learning and etc) to pFedACnnL for easy machine learning tasks can cause an over-fitting problem. Last but not least, pFedACnnL achieves the SOTA performance compared to the first- and zeroth-order PFL frameworks (i.e. pFedMe and pFedZO) in most cases while it achieves a competitive performance with pFedMe (96.23% versus 96.92%) in the non-convex setting of the MNIST dataset.
6 Conclusions
In this article, we have proposed FedACnnL to enable computation-efficient DNN training in AFL and proposed pFedACnnL based on it to mitigate the heterogeneous data distribution problem in AFL. Specifically, FedACnnL is the federated implementation of ACnnL with an introduction of the adaptive batch size algorithm to boost the training procedure. Meanwhile, pFedACnnL incorporates a gradient-free meta-learning method to generate a tailored model for each client’s data distribution. Experimentally, FedACnnL reduces the training time by $83\%\sim 99\%$ compared with the conventional first- and zeroth-order FL frameworks while pFedACnnL achieves the SOTA model performance in most cases of convex and non-convex settings.